OpenStudy (midhun.madhu1987):

A boat goes 40 km upstream in 8 hrs and 36 km downstream in 6 hrs. What is the speed of the boat in still water?

OpenStudy (perl):

so we can make a picture

OpenStudy (midhun.madhu1987):

Sure

OpenStudy (perl):

Let x = speed of boat in still water y = current of water Using distance = rate * time 40 = (x - y) * 8 36 = (x + y ) *6

OpenStudy (midhun.madhu1987):

Yeah... Thats what I wanted... so, they become: x - y = 5 x + y = 6 and solving them x = 5.5 and y = 0.5 Thanks a lot @perl